OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5). Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (anonymous):

Slope = (-4-5)/(-1-2) = -9/-3 = 3. So, y= 3x + b When x = 2 then y=5 : 5 = 3(2) + b => 5 = 6 + b b = -1. Hence equation of line is y = 3x - 1. This is not matter that what point we take at the end because slope is based on both the points.